OpenStudy (anonymous):

one sheepherder said to the other sheepherder give me 10 of your sheep to have twice as much as you. the other sheepherder answer him give me 10 of yours to have equal sheep. how many sheep dose each one have

OpenStudy (anonymous):

let the number of sheep owned by first sheep herder as A and the 2nd one is B. first sheepherder need another 10sheeps to hav twice of the second have.so the equation will be: A+10=2B now the second sheepherder need 10 more sheep to hav equal number.it can be written as: A+10=B u know how to solve the equations to get the answer?

OpenStudy (anonymous):

ok thanks thats a start do i 10/2= 5 so will B = 5

OpenStudy (anonymous):

sorry: for second equation, it should be B+10=A. NOT A+10=B

OpenStudy (anonymous):

ok no problem..

OpenStudy (anonymous):

what the answer u got?

OpenStudy (anonymous):

15 and 5

OpenStudy (anonymous):

A=15 and B= 5

OpenStudy (anonymous):

u should get A=30 and B=20

OpenStudy (anonymous):

o boy i suck... let me try it again

OpenStudy (anonymous):

well. im suck to.lol. let me start all over again

OpenStudy (anonymous):

the equations should be: 1) A+10= 2(B-10) 2) B+10= A-10

OpenStudy (anonymous):

lol aww but if A give B 10 it wil not be = because then A will have 20 in B will have 30 am im getting to into it huh

OpenStudy (anonymous):

for 1st equation, when farmer B give 10 of his sheeps to A, he will have less 10 sheeps and A have twice of his new amount. thats why A+10= 2(B-10). u understand this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but he say give me 10 so he can have = do he mean = as 30=30 or = as the same as u before u gave me 10

OpenStudy (anonymous):

for 2nd equation. they will have equal if A give 10 of his sheeps to B. so B hav additional 10sheeps and A hav less 10sheep. thats why: B+10 = A-10

OpenStudy (anonymous):

thats why u have equations: 1) A+10= 2(B-10) 2) B+10= A-10

OpenStudy (anonymous):

u got it kisha?

OpenStudy (anonymous):

ok i see i get it thanks well i did this other one on my own can u check it out for me 2 work are working together and their payment will b 10 $ one them work harder then the other and he will receive a 1$ more how much will each one be receiving the answer i got was A=6 B=4

OpenStudy (anonymous):

equations u should have: 1) A+B=10 2) A-B=1

OpenStudy (anonymous):

from (2); A-B=1 ---> A=1+B (3) substitute (3) into (1) yields; (1+B)+B =10 ---> B=(10-1)/2 = 4.50

OpenStudy (anonymous):

so A wil get $5.50.

OpenStudy (anonymous):

i hate math i have a big quiz coming up and im trap on problem like these its so hard to come up with a equation

OpenStudy (anonymous):

lol kisha. practice will make u better. u just have to learn to understand the situation. good luck for ur upcoming quiz. :)

OpenStudy (anonymous):

thanks and thank youuuuuuuuuuuuuuuuuuuu sooooooooooooooooo much words can explain